# Dialectal and Low Resource Machine Translation for Aromanian

### v1

@@ 9 Limitations
Aromanian is a very low resource language and is not standardized, having several varieties. Due to this, translation quality is experimental and oftentimes sub-par what a user might expect from a translation system for medium to high resource languages. This model is biased to produce the Aromanian variety spoken mostly in Romania. In addition, translation errors may occur where certain information is missed or altered. The words generated by the model might not be representative for speakers across all the regions and varieties where Aromanian is spoken.
Last but not least, this work mentions already a limitation in terms of lack of availablity to carry an in-depth speaker-centric human evaluation of the machine translation system. This is an issue we are looking into for an extended version of this work.
@@ 10 Ethical Considerations
We were able to collect and compile parallel translated texts from multiple sources. Note that for each text source, we have received explicit permission to use the text, the work is licensed in such a way that we are legally allowed to utilize and release it, or our scope falls under the fair use case.
11 Acknowledgements
This work would have not been possible without support from the Aromanian community, who helped us with the data collection processs, system testing, and gave us invaluable advice. We also acknowledge the support of the Human Language Technologies Research Center (University of Bucharest), for model training compute power, and hosting our online translation system.
Appendix A Aromanian-Romanian sample pairs per source
@@
Shi vasiljelu Ptolometi muri dupã trei dzali, shi-ashchirladzlji cari eara pit tsitãts lj-vatamara oaminjlji di-a loclui.
@@
| Radio |
Deputatlu USR Cristian Seidler spusi cã proiectul di nom ari un impact multu modestu ti hãrgili bugetare cu pensiile spetsiali.
Timishoara fu singurã dinintea a represiuniljei trǎ ma multu di 5 dzãle di foc shi sãndze. |
Deputatul USR Cristian Seidler a afirmat că proiectul de lege are un impact foarte modest asupra cheltuielilor bugetare cu pensiile speciale.
Timișoara a fost singură în fața represiunii pentru cel puțin 5 zile de foc și sânge. |
| Aromanian Poetry |
@@
Our final data release contains an additional k dictionary pairs that we do not use to train the models. When training our initial machine translation models, we did not observe any improvements when using word-aligned dictionary entries. The entries are extracted from Papahagi’s Aromanian dictionary Papahagi (1963).
Appendix B Corpus statistics per source
@@
| language | source | words | unique words | words/sentence |
| Aromanian | Bible | 493018 | 69780 | 16.15 |
| Lyrics T. | 16039 | 7586 | 3.98 | |
| Small corpus | 508090 | 74419 | 27.1 | |
| Phrases | 40941 | 9972 | 4.78 | |
| Avdhela | 7839 | 3567 | 3.48 | |
| Radio | 41036 | 12708 | 12.15 | |
| Aromanian Poetry | 36649 | 10433 | 11.69 | |
| Tom Sawyer | 23177 | 10325 | 13.13 | |
| Divine Comedy | 1269 | 747 | 4.83 | |
| Modern poetry | 11897 | 6336 | 5.3 | |
| Tao Te Ching | 5653 | 1886 | 2.62 | |
| The little prince | 3619 | 1625 | 5.81 | |
| Writings Collection | 17893 | 6646 | 12.81 | |
| Romanian | Bible | 580288 | 42487 | 19.0 |
| Lyrics T. | 16923 | 7022 | 4.2 | |
| Small corpus | 494098 | 53549 | 26.36 | |
| Phrases | 43590 | 7484 | 5.08 | |
| Avdhela | 8186 | 3330 | 3.63 | |
| Radio | 47588 | 12291 | 14.09 | |
| Aromanian Poetry | 38719 | 11126 | 12.35 | |
| Tom Sawyer | 24665 | 9296 | 13.97 | |
| Divine Comedy | 1397 | 744 | 5.31 | |
| Modern poetry | 14146 | 6412 | 6.31 | |
| Tao Te Ching | 4877 | 1740 | 2.26 | |
| The little prince | 3977 | 1717 | 6.38 | |
| Writings Collection | 19328 | 6398 | 13.84 | |
| English | Bible | 622430 | 28266 | 20.38 |
| Lyrics T. | 20277 | 5067 | 5.03 | |
| Small corpus | 518118 | 34576 | 27.64 | |
| Phrases | 49025 | 4950 | 5.72 | |
| Avdhela | 9466 | 2671 | 4.2 | |
| Radio | 53286 | 8348 | 15.78 | |
| Aromanian Poetry | 41069 | 7692 | 13.1 | |
| Tom Sawyer | 29411 | 6491 | 16.67 | |
| Divine Comedy | 1457 | 675 | 5.54 | |
| Modern poetry | 16959 | 4892 | 7.56 | |
| Tao Te Ching | 5089 | 1597 | 2.36 | |
| The little prince | 4258 | 1380 | 6.83 | |
| Writings Collection | 20998 | 4795 | 15.03 |
Appendix C LaBSE training hypermateres
@@
Appendix D NLLB training hypermateres
@@ E.2 Prompt format
We format the prompts of the LLaMA 3.1 8B Instruct and RoLLaMA 3 8B Instruct models with the standard LLaMA 3.1 Instruct prompt format Dubey et al. (2024). For the Qwen and TowerInstruct models, we use their specific instruction prompt template Yang et al. (2024); Alves et al. (2024).
@@
Appendix F Automatic evaluation results
| model | ron-¿rup | rup-¿ron | eng-¿rup | rup-¿eng | eng-¿ron | ron-¿eng |
@@ Appendix F Automatic evaluation results
| Qwen2 7B Instruct | 15.2 | 28.8 | 10.3 | 32.4 | 35.4 | 62.9 |
| RoLLaMA | 14.7 | 29.7 | 11.0 | 32.7 | 38.1 | 64.4 |
| TowerInstruct v0.2 | 16.2 | 29.4 | 7.3 | 23.6 | 27.1 | 62.0 |
| LLaMA 3.1 8B Instruct | 15.8 | 30.2 | 10.7 | 33.7 | 38.6 | 65.1 |
| NLLB 1.3B | 16.2 | 31.9 | 12.8 | 35.6 | 41.4 | 63.0 |
| NLLB 600M | 17.0 | 30.9 | 13.4 | 34.9 | 39.1 | 62.2 |
| model | ron-¿rup | rup-¿ron | eng-¿rup | rup-¿eng | eng-¿ron | ron-¿eng |

### v2

@@ 9 Limitations
Aromanian is a very low-resource language and is not standardized, having several varieties. As a result, translation quality is experimental and often times sub-par to what a user might expect from a translation system for medium- to high-resource languages. This model is biased to favour some Aromanian varieties more than others. Translation errors may occur when certain information is missing or altered. There is also concern about the model’s bias, as the Bible is a dominant source in our corpus. Last but not least, this work already mentions a limitation in terms of lack of availability to carry an in-depth speaker-centric human evaluation of the MT system. This is an issue that we are looking into in future work. We release the corpus under the Creative Commons Attribution-NonCommercial 4.0 International License 111111https://creativecommons.org/licenses/by-nc/4.0/.
@@ 10 Ethical Considerations
We were able to collect and compile parallel translated texts from multiple sources. Note that for each text source we have received permission to use the text, verified appropriate licensing agreements, or confirmed fair-use applicability for academic research purposes. The Google Translate API calls cost around $ in credits and the OpenAI API calls cost around $. Model training was performed on cloud-based servers (A100 and RTX 4090 GPUs) rented by the hour, incurring a total cost of approximately $.
Acknowledgments
This work would not have been possible without the support of the Aromanian Community from Romania (CARo), who helped us with the data collection processes, system testing, and gave us invaluable advice. We specifically acknowledge Florentina Costea, Cristian Jiru, and Nicolae Todică. Regarding data collection, we are grateful to Kira Mantsu and George Vrana for all their help. The human evaluation process would not have been possible without the generous help of Georgiana Stavrositu, Sebastian Florea, and Zoia Dragomir. The machine translation system is hosted with the generous support of the University of Bucharest https://arotranslate.unibuc.ro.
This research is partially supported by the project “Romanian Hub for Artificial Intelligence - HRIA”, Smart Growth, Digitization and Financial Instruments Program, 2021-2027, MySMIS no. 334906 and by InstRead: Research Instruments for the Text Complexity, Simplification and Readability Assessment CNCS - UEFISCDI project number PN-IV-P2-2.1-TE-2023-2007.
Appendix A Aromanian-Romanian Sample Pairs
@@
Și vasiljelu Ptolometi muri după trei dzali, shi-ashchirladzlji cari eara pit tsitãts lj-vatamara oaminjlji di-a loclui.
@@
| Radio | Deputatlu USR Cristian Seidler spusi cã proiectul di nom ari un impact multu modestu ti hãrgili bugetare cu pensiile spetsiali. | Deputatul USR Cristian Seidler a afirmat că proiectul de lege are un impact foarte modest asupra cheltuielilor bugetare cu pensiile speciale. |
Aromanian
Poetry |
@@
| Source | Aromanian | Romanian |
| Bible |
Și vasiľelu Ptolometi muri dupâ trei dzali, și-așchirladzľi cari eara pit țităț ľ-vatamara oamińľi di -a loclui.
Ti-agiutarea-a ților far di mindueari-arauda unâ giudicata cu ițraľi, a tinirlui om ti-unâ cânușteari și bunâ mindueari. |
| Lyrics T. |
Pisti giuguri trecu anarga, lâi niori
nicuchirâ cu curunâ: |
Peste juguri trec încet, nori negri
Ar trimite-o el acasă, |
| MPC-rup |
Unâ caprâ mușcră și unâ ghesucănută pășteau di unâ parte, și chipurle lâ asunau: cing-cing.
Țachiľi giuca deavârliga, arsărea-tuț alepțâ, mușeaț, cu lâna lor ca mătasea, di-ľi yinea și-al Tegâ se-arsară și s-lu stringă-n brațâ picurarlu și s-lu bașe, ahtare harauâ lu-avea loatâ! |
| Phrases |
nu știu si mpartu paľi la doi yumari
misuru stealili |
fi prost
hoinări |
| Avdhela |
Di toamnă-ascuturată Ma,
s-ti caftu ńi-easti fricâ, |
Se scutură de toamnă,
Să te cer îmi este teamă, |
| Radio | Deputatlu USR Cristian Seidler spusi că proiectul di nom ari un impact multu modestu ti hârgili bugetare cu pensiile spețiali. | Deputatul USR Cristian Seidler a afirmat că proiectul de lege are un impact foarte modest asupra cheltuielilor bugetare cu pensiile speciale. |
Aromanian
Poetry |
Am un alt livend, di-Avdelă, alâvdatlu-ațel di NUȘI, te da ńeate cât ș-la-auși;
nâ șcâmbâ fârâ suflit i semn di tine, tutâ și salta siminată di tine, tu ubor… |
Tom
Sawyer |
Arucâ funea!
Mutri anvârliga, canâ… tu niheam di oarâ u avea cartea tru mânâ. |
Dă și parâma!
Divine
Comedy |
Și-agârșeaști-n frixea ți-ľi da dăgoarea
tu doiľi șerchi-mpriunaț trupeaști |
pârjolul să-și ia ceva, căci la copil ia seamă
în cei doi șerpi împreunați |
Modern
poetry |
mași suflitlu nu ari moarti
aclo iu suntu farurli apâryisiti. |
doar sufletul este fără moarte
acolo unde sunt farurile abandonate. |
Tao
Te Ching |
Poț s-ti tradz ditu minduirea a ta
Imńi dupâ ea ș-nu-ari bitisitâ. |
Poți să te dai înapoi din propria ta minte
urmează-l și nu are sfârșit. |
The Little
Prince |
Gioaca cor cathi gioi cu featili di-n hoara.
Eľ avdu daima mași alavdarli. |
Se duc să joace, joia, cu fetele din sat.
Vanitoșii nu aud niciodată decât laudele. |
| Writings |
Li-alidzea ayońia, unâ dupu altâ, ași cum beai unâ scafâ cu apâ!
Bânarâ deadunu tu vâryârie. |
Le citea repede, una după alta, așa cum ai bea un pahar cu apă!
Trăiseră împreună în Bulgaria. |
@@
Our final data release also contains k dictionary pairs that we do not use to train the models. When training our initial machine translation models, we did not observe any improvements when using word-aligned dictionary entries. The entries are extracted from Papahagi’s Aromanian dictionary Papahagi (1974).
Appendix B Corpus Statistics
@@
| language | source | words | unique words | words/sentence | type-token ratio |
| Aromanian | Bible | 493018 | 69780 | 16.15 | 0.14 |
| Lyrics T. | 16039 | 7586 | 3.98 | 0.47 | |
| MPC-rup | 508090 | 74419 | 27.1 | 0.14 | |
| Phrases | 40941 | 9972 | 4.78 | 0.24 | |
| Avdhela | 7839 | 3567 | 3.48 | 0.45 | |
| Radio | 41036 | 12708 | 12.15 | 0.30 | |
| Aromanian Poetry | 36649 | 10433 | 11.69 | 0.28 | |
| Tom Sawyer | 23177 | 10325 | 13.13 | 0.44 | |
| Divine Comedy | 1269 | 747 | 4.83 | 0.58 | |
| Modern poetry | 11897 | 6336 | 5.3 | 0.53 | |
| Tao Te Ching | 5653 | 1886 | 2.62 | 0.33 | |
| The little prince | 3619 | 1625 | 5.81 | 0.44 | |
| Writings Collection | 17893 | 6646 | 12.81 | 0.37 | |
| Romanian | Bible | 580288 | 42487 | 19.0 | 0.07 |
| Lyrics T. | 16923 | 7022 | 4.2 | 0.41 | |
| MPC-rup | 494098 | 53549 | 26.36 | 0.10 | |
| Phrases | 43590 | 7484 | 5.08 | 0.17 | |
| Avdhela | 8186 | 3330 | 3.63 | 0.40 | |
| Radio | 47588 | 12291 | 14.09 | 0.25 | |
| Aromanian Poetry | 38719 | 11126 | 12.35 | 0.28 | |
| Tom Sawyer | 24665 | 9296 | 13.97 | 0.37 | |
| Divine Comedy | 1397 | 744 | 5.31 | 0.53 | |
| Modern poetry | 14146 | 6412 | 6.31 | 0.45 | |
| Tao Te Ching | 4877 | 1740 | 2.26 | 0.35 | |
| The little prince | 3977 | 1717 | 6.38 | 0.43 | |
| Writings Collection | 19328 | 6398 | 13.84 | 0.33 | |
| English | Bible | 622430 | 28266 | 20.38 | 0.04 |
| Lyrics T. | 20277 | 5067 | 5.03 | 0.24 | |
| MPC-rup | 518118 | 34576 | 27.64 | 0.06 | |
| Phrases | 49025 | 4950 | 5.72 | 0.10 | |
| Avdhela | 9466 | 2671 | 4.2 | 0.28 | |
| Radio | 53286 | 8348 | 15.78 | 0.15 | |
| Aromanian Poetry | 41069 | 7692 | 13.1 | 0.18 | |
| Tom Sawyer | 29411 | 6491 | 16.67 | 0.22 | |
| Divine Comedy | 1457 | 675 | 5.54 | 0.46 | |
| Modern poetry | 16959 | 4892 | 7.56 | 0.28 | |
| Tao Te Ching | 5089 | 1597 | 2.36 | 0.31 | |
| The little prince | 4258 | 1380 | 6.83 | 0.32 | |
| Writings Collection | 20998 | 4795 | 15.03 | 0.22 |
Appendix C LaBSE Training Hyper-paramateres
@@
Appendix D NLLB Training Hyper-parameters
@@ E.2 Prompt Format
We format the prompts of the LLaMA 3.1 8B Instruct and RoLLaMA 3 8B Instruct models using the standard LLaMA 3.1 Instruct prompt format Grattafiori et al. (2024). For the Qwen and TowerInstruct models, we use their specific instruction prompt template Yang et al. (2024); Alves et al. (2024).
@@
Appendix F Automatic Evaluation Results - test Set
| model | ronrup | rupron | engrup | rupeng | engron | roneng |
@@ Appendix F Automatic Evaluation Results - test Set
| Qwen2 7B Instruct | 15.2 | 28.8 | 10.4 | 32.4 | 35.4 | 63.0 |
| TowerInstruct v0.2 | 16.3 | 29.5 | 7.3 | 23.7 | 27.1 | 62.1 |
| LLaMA 3.1 8B Instruct | 15.9 | 30.3 | 10.8 | 33.7 | 38.6 | 65.1 |
| RoLLaMA | 14.8 | 29.7 | 11.1 | 32.8 | 38.1 | 64.4 |
| NLLB 1.3B | 16.3 | 31.9 | 12.8 | 35.7 | 41.4 | 63.0 |
| NLLB 600M | 17.0 | 30.9 | 13.5 | 35.0 | 39.1 | 62.2 |
| model | ronrup | rupron | engrup | rupeng | engron | roneng |
@@
Appendix G Automatic Evaluation Results - dev Set
| model | ronrup | rupron | engrup | rupeng | engron | roneng |
| Qwen2 7B Instruct | 31.9 | 50.7 | 23.9 | 51.7 | 49.8 | 67.8 |
| TowerInstruct v0.2 | 34.5 | 52.6 | 26.4 | 53.1 | 52.1 | 70.5 |
| LLaMA 3.1 8B Instruct | 34.8 | 53.6 | 26.6 | 52.4 | 52.4 | 69.5 |
| RoLLaMA | 34.4 | 53.0 | 26.3 | 51.4 | 52.4 | 68.9 |
| NLLB 1.3B | 33.7 | 53.2 | 26.2 | 52.7 | 52.1 | 68.5 |
| NLLB 600M | 33.6 | 52.0 | 25.6 | 52.2 | 50.5 | 67.3 |
| model | ronrup | rupron | engrup | rupeng | engron | roneng |
| Qwen2 7B Instruct | 61.9 | 68.1 | 55.6 | 67.6 | 70.1 | 80.3 |
| TowerInstruct v0.2 | 63.3 | 69.1 | 56.9 | 68.8 | 71.9 | 82.1 |
| LLaMA 3.1 8B Instruct | 63.7 | 69.8 | 57.2 | 68.5 | 72.0 | 81.6 |
| RoLLaMA | 63.7 | 69.4 | 57.2 | 68.1 | 71.9 | 81.1 |
| NLLB 1.3B | 63.7 | 69.9 | 57.7 | 68.4 | 72.2 | 80.8 |
| NLLB 600M | 63.6 | 69.2 | 57.4 | 67.9 | 71.1 | 80.0 |
Compared to the out-of-domain test set, the evaluation scores on the dev set are considerably higher and the differences between LLMs and NLLB-type models is smaller.
ChrF signature is nrefs:1|case:mixed|eff:yes|nc:6|nw:0|space:no|version:2.4.3 and BLEU signature is nrefs:1|case:mixed|eff:no|tok:13a|smooth:exp|version:2.4.3.